# AdaCQR: Enhancing Query Reformulation for Conversational Search via Sparse and Dense Retrieval Alignment

### v1

@@ Appendix A Discussion
A.1 Effectiveness of Prompt Setting
@@
A.2 Generalization on Out-Of-Domain (OOD) Dataset
Sparse Dense
Model MRR NDCG MRR NDCG
ConvGQR(ID) 12.4 10.7 25.6 24.3
IterCQR(OOD) 13.7 12.2 17.8 16.4
AdaCQR(OOD) 14.0 12.6 20.1 18.6
Table 6: Performance of AdaCQR on out-of-distribution dataset. We use the AdaCQR model trained on QReCC and test on TopiOCQA dataset.
To measure the generalization performance of AdaCQR, we trained the model on the QReCC dataset and evaluated it on the TopiOCQA dataset, with the results presented in Table 6. As indicated by the results in Table 6, AdaCQR demonstrates superior generalization performance, outperforming IterCQR in both sparse and dense Retrieval, and surpassing the in-domain model ConvCQR in sparse retrieval.
@@ Appendix B Experimental Details
B.1 Datasets Details
The QReCC dataset comprises 14K conversations with 80K question-answer pairs, and we aim to retrieve the gold passage from a collection containing 54M passages. Conversely, the TopiOCQA dataset includes 3.9K topic-switching conversations with 51K question-answer pairs, where the passage collection is sourced from Wikipedia and contains about 20M passages. Notably, a few examples from the QReCC and TopiOCQA training sets were randomly partitioned to create respective validation sets. The datasets details are described in Table 7.
@@ B.1 Datasets Details
B.2 Evaluation Metrics
@@ C.1 AdaCQR Details
AdaCQR is trained for $10$ epochs in Stage 1 with a learning rate set to 2e-5 and 8 epochs in Stage 2 with a learning rate adjusted to 5e-6. Both stages incorporate linear learning rate schedulers with a warm-up ratio of $0.1$.
@@ C.2 Retrieval Systems Details
We implement the retrieval systems using Faiss Johnson et al. (2019) and Pyserini Lin et al. (2021a). For BM25, as in previous work Mo et al. (2023a); Jang et al. (2023); Yoon et al. (2024), we set $k_{1}=0.82$, $b=0.68$ in QReCC, and $k_{1}=0.9$, $b=0.4$ in TopiOCQA. The $k_{1}$ controls the non-linear term frequency normalization and $b$ is the scale of the inverse document frequency. For ANCE [^6] https://huggingface.co/sentence-transformers/msmarco-roberta-base-ance-firstp, the maximum token length is set to 128 tokens for reformulation query and 384 tokens for passage.
@@ Appendix D ChatGPT Annotation Details
For superior reformulation labels for AdaCQR, we utilize the top-3 most challenging demonstrations (i.e., $m=3$) for the QReCC dataset and the top-5 most challenging demonstrations (i.e., $m=5$) for the TopiOCQA dataset, i.e., few-shots. The prompts to annotate the QReCC dataset and the TopiOCQA dataset are shown in Table 8 and Table 9, respectively.
To encourage a more deterministic output, we set the temperature to $0.1$ and seed is set to $42$ for reproductivity. The total consumption to annotate QReCC and TopiOCQA datasets for initial and superior reformulation labels is about 151M tokens, which cost about 120\$.
@@ Appendix E Case Study
Table 12 demonstrates a case where AdaCQR successfully retrieved the gold passage through query rewriting, whereas human rewrites failed, showing the superiority of AdaCQR over human rewrites. After being written by AdaCQR, the query is decontextualized, resulting in overlaps while concurrently offering more specific information. This enhanced specificity aids the retriever toward the most relevant passages effectively. Additionally, in Tables 13 and 14, we also show examples of how the AdaCQR and AdaCQR with Expansion models successfully retrieved the gold passage.

### v3

@@ Appendix A Discussion
A.1 Advantage of AdaCQR in Alignment
Previous works like CONQRR Wu et al. (2022) and IterCQR Jang et al. (2023) explored the concept of alignment between the reformulation model and the retriever. CONQRR leverages reinforcement learning to maximize the reward of the sampled rewrite $q_{s}$ and the baseline rewrite $q$. IterCQR uses Minimum Bayes Risk (MBR) and Top-1 Selection to achieve the alignment between the reformulation model and dense retriever.
Compared with previous works, the main advantage of AdaCQR is that it no longer requires an explicit reward model during training; the language model serves simultaneously as both the generation model and the evaluation model (the evaluation model of CONQRR is BM25 retrieval system, and the evaluation model of IterCQR is a dense passage encoder). By decoupling the model training and candidates ranking (i.e., the candidates ranking is an offline process), our method achieves improved training efficiency.
Apart from this, AdaCQR achieves alignment based on contrastive learning, which is easier to implement, converges faster, and is more stable during training compared to the reinforcement learning(RL) of CONQRR and the Minimum Bayes Risk(MBR) of IterCQR  Liu et al. (2022). Both CONQRR and IterCQR employ special techniques like min-max normalization to ensure model convergence, which increases the training complexity (Wu et al., 2022; Jang et al., 2023). Additionally, IterCQR requires two stages, MBR and Top-1 Selection, to achieve alignment in 7-10 epochs Jang et al. (2023), whereas AdaCQR achieves alignment in just 3-5 epochs.
A.2 Effectiveness of Prompt Setting
@@
A.3 Effect of the Multi-Task Loss
QReCC
Coefficient($\gamma$) MRR NDCG R@10 R@100
0 43.3 41.0 62.8 88.5
0.1 45.3 42.7 65.2 90.2
1 48.8 46.1 68.7 91.2
10 50.2 47.7 68.8 89.0
100 52.4 49.9 70.9 91.0
1000 49.4 46.7 68.6 90.7
$\infty$ 44.5 41.8 65.5 90.9
A.4 Robustness to Topic Shifts in Conversation
Topic-Concentrated Topic-Shifted
Model MRR R@10 MRR R@10
T5QR 35.2 54.4 25.2 45.1
CONQRR 41.9 63.1 25.2 45.9
IterCQR 54.4 72.4 24.9 49.7
Human Rewrite 44.0 66.7 31.8 56.7
AdaCQR 66.0 82.4 34.1 58.3
Table 8: Performance of AdaCQR on topic-concentrated and topic-shifted samples on QReCC, MRR and R@10 are reported. The result is reported on the BM25 Retrieval System.
@@ Appendix B Experimental Details
B.1 Baseline Details
In our study, we compare AdaCQR with the following representative baselines in the CQR task:
  • T5QR Lin et al. (2020) is a vanilla baseline that utilizes the T5-base Raffel et al. (2020) model to perform CQR tasks.
  • CONQRR Wu et al. (2022) aligns the T5-base reformulation model with retrievers through direct optimization using reinforcement learning.
  • ConvGQR  Mo et al. (2023a) enhances retrieval performance by employing two fine-tuned T5-base models, one for query reformulation and the other for query expansion.
  • EDIRCS  Mao et al. (2023a) leverages the incorporation of non-autoregressive text-selecting techniques and autoregressive tokens generation to generate reformulation queries effectively on a fine-tuned T5-base model.
  • LLM-Aided  Ye et al. (2023) employs ChatGPT  OpenAI (2022) to conduct query reformulation via a “rewrite-then-edit” prompting strategy.
  • IterCQR  Jang et al. (2023) achieves the alignment of the T5-base reformulation model and dense retriever by minimizing Bayes Risk based on the semantic similarity between the query and the gold passage.
  • RetPO Yoon et al. (2024) utilizes large language models to generate multiple reformulations through multi-perspective prompting, creates binarized comparisons based on retriever feedback, and optimizes LLaMA2-7B Touvron et al. (2023) using direct preference optimization (DPO) Rafailov et al. (2023).
  • LLM4CS Mao et al. (2023b) is the state-of-the-art conversational query reformulation baseline, utilizing LLM prompting techniques to generate multiple reformulated queries and aggregate their embeddings. We adopt rewriting prompting (REW) and rewriting-and-response (RAR) as baselines for query reformulation and query reformulation with expansion, respectively. Our implementation of REW and RAR follows standard settings (i.e., generating five queries and aggregating them) but excludes chain-of-thought (CoT) content due to the extensive human annotation required, which is not part of our baselines and impractical in real-world scenarios. It is important to note that aggregation techniques are applicable only in dense retrieval; therefore, for BM25 retrieval, we implement RAR by concatenating the query and response with the top-1 generation result.
B.2 Datasets Details
@@
The QReCC Anantha et al. (2021) dataset comprises 14K conversations with 80K question-answer pairs, and we aim to retrieve the gold passage from a collection containing 54M passages. Conversely, the TopiOCQA Adlakha et al. (2022) dataset includes 3.9K topic-switching conversations with 51K question-answer pairs, where the passage collection is sourced from Wikipedia and contains about 20M passages. Notably, a few examples from the QReCC and TopiOCQA training sets were randomly partitioned to create respective validation sets. The details of QReCC and TopiOCQA are described in Table 9.
TREC CAsT 2019, 2020, and 2021 Dalton et al. (2020, 2021, 2022) are datasets known for their complexity and challenges in conversational search with a zero-shot setting. The statistics of CAsT 19-21 are shown in Table 10.
CAsT-19 CAsT-20 CAsT-21
# Dialogues 50 25 26
# Turns 479 208 239
# Collections 38M 38M 40M
Table 10: The statistics of CAsT 2019, 2020, and 2021 datasets.
B.3 Evaluation Metrics
@@ C.1 AdaCQR Details
AdaCQR is trained for $10$ epochs in Stage 1 with a learning rate set to 2e-5 and 8 epochs in Stage 2 with a learning rate adjusted to 5e-6. Both stages incorporate linear learning rate schedulers with a warm-up ratio of $0.1$. We employ grid search for hyperparameter selection.
@@ C.2 Retrieval Systems Details
We implement the retrieval systems using Faiss Johnson et al. (2019) and Pyserini Lin et al. (2021a). For BM25, as in previous work Mo et al. (2023a); Jang et al. (2023); Yoon et al. (2024), we set $k_{1}=0.82$, $b=0.68$ in QReCC, and $k_{1}=0.9$, $b=0.4$ in TopiOCQA. The $k_{1}$ controls the non-linear term frequency normalization, and $b$ is the scale of the inverse document frequency. For ANCE [^6] https://huggingface.co/sentence-transformers/msmarco-roberta-base-ance-firstp, the maximum token length is set to 128 tokens for reformulation query and 384 tokens for passage.
@@ Appendix D ChatGPT Annotation Details
For superior reformulation labels for AdaCQR, we utilize the top-3 most challenging demonstrations (i.e., $m=3$) for the QReCC dataset and the top-5 most challenging demonstrations (i.e., $m=5$) for the TopiOCQA dataset, i.e., few-shots. The final prompt is formed $\mathcal{D}=\mathcal{I}\ ||T_{1}||\cdots||T_{m}$, where ||||| | donates concatenation. The detailed prompts to annotate the QReCC dataset and the TopiOCQA dataset are shown in Table 11 and Table 12, respectively.
To encourage a more deterministic output, we set the temperature to $0.1$, and the seed is set to $42$ for reproductivity. The total consumption to annotate QReCC and TopiOCQA datasets for initial and superior reformulation labels is about 151M tokens, which cost about 120\$.
@@ Appendix E Case Study
Table 15 demonstrates a case where AdaCQR successfully retrieved the gold passage through query rewriting, whereas human rewrites failed, showing the superiority of AdaCQR over human rewrites. After being written by AdaCQR, the query is decontextualized, resulting in overlaps while concurrently offering more specific information. This enhanced specificity aids the retriever toward the most relevant passages effectively. Additionally, in Tables 16 and 17, we show examples of how the AdaCQR and AdaCQR with Expansion models successfully retrieved the gold passage. We also provide a failure case, as illustrated in Table 18.
@@ Appendix F Query Expansion Details
Conversation:
Q1: How can you tell if someone is suffering from depression?
A1: You may be depressed if, for more than two weeks, you’ve felt sad, down or miserable most of the time, or have lost interest or pleasure in usual activities, and have also experienced several of the signs and symptoms across at least three of the categories below.
Q2: What are common types?
A2: he four most common types of depression are major depression, persistent depressive disorder(formerly known as dysthymia), bipolar disorder, and seasonal affective disorder.
Q3: What causes it?
A3: Rather, there are many possible causes of depression, including faulty mood regulation by the brain, genetic vulnerability, stressful life events,
Q4: What is the role of brain chemicals?
A4: A chemical imbalance in the brain is said to occur when there’s either too much or too little of certain chemicals, called neurotransmitters, in the brain. Neurotransmitters are natural chemicals that help facilitate communication between your nerve cells.
Q5: Does a lack of sunlight cause it?
A5: If you’re not careful, a lack of sunlight can actually lead to a form of clinical depression.
Original Query: How can you treat SAD? (rank: 2)
Human Rewrite: How can you treat SAD? (rank: 2)
AdaCQR(Ours): How can SAD be treated if a lack of sunlight in the brain can actually lead to depression? (rank: Not Found)
Table 18: Failure case on QReCC (id: 57_6) when using BM25.